OpenStudy (anonymous):

verify the identity cot (theta-pi/2)=-tan(theta)

OpenStudy (anonymous):

try writing that left side in terms of cosine and sine.... what u got?

OpenStudy (anonymous):

after you do that, use the sum and difference rule for sine and cosine.... you should be able to simplify easily from there....

OpenStudy (anonymous):

hold on

OpenStudy (abb0t):

\(\cot(x) = \frac{ \cos(x) }{ \sin(x) }\) and \(\tan(x) = \frac{ \sin(x) }{ \cos(x) }\)

OpenStudy (anonymous):

ok can u show me how to get that

OpenStudy (anonymous):

as @abb0t stated, cotx = cosx/sinx.... so \(\large cot(\theta - \pi/2)=\frac{cos(\theta-\pi/2)}{sin(\theta-\pi/2)}\)

OpenStudy (anonymous):

now, use the difference rules for sine and cosine...

OpenStudy (anonymous):

here.... check out page 2: http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

ok its sin(x)cos(pi/2) - cos(x)sin(pi/2)

OpenStudy (anonymous):

yes... that's for the denominator.... how 'bout the numerator?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

cos(x)cos(pi/2) - sin(x)sin(pi/2)

OpenStudy (anonymous):

very nice.... now, cos(pi/2) = ??? sin(pi/2) = ???

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

oh wait.... cos(x - pi/2) = cosx cos(pi/2) + sinx sin(pi/2) this is correct.... make sure you make the change....

OpenStudy (anonymous):

ok i have to change the - to a +

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hang on.... afk...

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

afk: away from keyboard.... but i'm back now...

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

i have a another problem i need help with